# Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer

### v1

@@ III.2.2 Classical resource details
For the [4Fe-4S] experiments, the eigenstate solver can be distributed across a number of classical nodes. The workflow for [4Fe-4S] is run on the supercomputer Fugaku. The supercomputer Fugaku has a total of 158,976 nodes with Armv8.2-A SVE 512 bit architecture. Each node has 82 cores and 32 GiB of memory. The largest calculation in this work uses 6400 nodes of the Fugaku supercomputer running in parallel.
IV Experimental results
@@
IV.1 Visualization of the wave function amplitudes
@@ IV.2 Scaling and runtime
Fig. S11 shows the potential energy surfaces obtained by the estimator with different values of $d$, the number of determinants used for the subspace expansion and diagonalization. Increasing the value of $d$ improves the results both qualitatively and quantitatively. For the smaller values of $d$, the potential energy surface shows unphysical oscillations when the bond length is large. For larger values of $d$, not only the energy values are lower, but also the oscillations in the energy as a function of the bond length are significantly decreased in amplitude. Improvement of the ground state approximation with an increasing amount of computational resources is also observed in the energy-variance analyses of the main text (Fig. 4), and the supplementary materials (Figs. S20 and S21).
@@
IV.3 Effect of orbital optimizations in the dissociation of N2 (cc-pVDZ)
@@
IV.4 Orbital occupation numbers for the low-energy spectrum of [2Fe-2S]
@@
IV.5 Quantum signal in the experiments
@@
The third set of tests involves comparing the performance of the HPC quantum estimator with configuration recovery run on configurations obtained from the quantum device and samples obtained from the uniform distribution. We consider the uniform distribution over all possible configurations in the Fock space, as well as the uniform distribution over configurations with the correct particle number, as in Fig. 4 (c) in the main text  [^1] When considering samples from the uniform distribution over electronic configurations on the correct particle sector, the configuration recovery method cannot be applied since it only recovers configurations with the wrong particle number. Therefore, this scenario is equivalent to running QSCI (without configuration recovery) on those samples.. The results for these last two sets of tests are shown in the coming subsections for the 40-, 52- and 72-qubit experiments.
IV.5.1 N2 experiments

### v2

@@ III.2.2 Classical resource details
For the [4Fe-4S] experiments, the eigenstate solver can be distributed across a number of classical nodes. The workflow for [4Fe-4S] is run on the supercomputer Fugaku. The supercomputer Fugaku has a total of 158,976 nodes with Armv8.2-A SVE 512 bit architecture. Each node has 48 compute cores and 32 GiB of memory. The largest calculation in this work uses 6400 nodes of the Fugaku supercomputer running in parallel.
IV Quantum circuits that can sample more efficiently than classical Selected Configuration heuristics
In this section, we show that there are LUCJ circuits that can sample electronic configurations more efficiently than some classes of classical heuristics, yielding equal or better energies than classical heuristics at fixed subspace dimensions.
For our study we consider a particular flavor of SCI, Heat bath Configuration Interaction (HCI) [34], and the [2Fe-2S] cluster, representative of strongly correlated species (20 orbitals $\rightarrow$ 40+ qubits). The circuit under consideration is the same LUCJ circuit that was used for the hardware experiments shown in the main text, with the same Heavy-Hex connectivity.
We first run HCI with different values of the HCI selection cutoff $\epsilon_{1}$ [81], as shown in Fig. S10. We observe that the larger and more restrictive values of $\epsilon_{1}$ do not allow HCI to reach the DMRG energy reference. Instead, it converges to the first excited state (S1) energy. Decreasing the value of $\epsilon_{1}$ allows the subspace dimension to become larger and improve the energy beyond that first excited state, at the cost of a significant increase in subspace dimension, which results in a higher computational cost. We note that for $\epsilon_{1}>8\cdot 10^{-5}$ it takes two or more HCI iterations to leave the first excited state. From the largest HCI wave function (last point for $\epsilon_{1}=5\cdot 10^{-5}$), we repeatedly remove the electronic configurations whose wave function amplitude is below a given threshold and compute the energy in the resulting subspace, providing a collection of energy-dimension pairs that we label as “upper bound to optimal” in Fig. S10. The removal of tens of millions of configurations (resulting in significantly smaller subspace dimensions) does not significantly deteriorate the quality of the ground-state approximation. Eventually, when the subspace dimension becomes smaller than $50\cdot 10^{6}$, the deterioration of the energy becomes more notable. These results indicates that HCI does not perform an optimal search of the relevant electronic configurations. Consequently, much smaller subspaces exist that yield comparable energy values. In this particular molecule, HCI needs to explore and perform diagonalizations in subspaces much larger than the optimal search.
After assessing that there exist subspaces of smaller dimension that yield equivalent energy estimations to HCI, we optimize an LUCJ circuit to produce samples in the identified subspaces with high probability. The optimal parameters are found in a two step optimization workflow. First, we optimize the Kullback–Leibler divergence between samples drawn from the LUCJ circuit and the amplitudes of 58 M-dimensional ground-state wave function estimation (green arrow in Fig. S10). Then, the circuit parameters are further fine-tuned to minimize the energy using a differential-evolution strategy.
With the samples produced by the optimized LUCJ circuit, we proceed to evaluate the energy of resulting subspaces of varying dimensionality, including the 58 M-dimensional subspace that the circuit was optimized for. We observe good agreement between the LUCJ subspaces and those that are an upper bound to the optimal ones. It is worth noting that despite being optimized for a 58 M-dimensional subspace, the resulting circuit produces subspaces of outstanding quality of dimensionalities both much smaller and larger. This result highlights the high ability of the LUCJ circuit to explore the space of electronic configurations.
Figure S10: Comparison between HCI and optimized LUCJ circuits in the efficiency of exploration of relevant electronic configurations to describe the ground state of [2Fe-2S]. The panels show the energy estimation as a function of subspace dimension of the diagonalization required to reach the corresponding energy. The dots connected by arrows show the trajectory of HCI in the energy-dimension plane. Each dot is labelled by the iteration it corresponds to. Different panels correspond to HCI calculations carried out with different values of $\epsilon_{1}$, as indicated in the legend. Brown triangles show energies and subspace dimensions obtained after truncating the most accurate HCI wave function obtained by repeatedly removing the electronic configurations of lowest amplitude. The green triangles show the energies for different subspace dimensions defined by samples obtained from an optimized LUCJ circuit. The green arrows show the subspace dimension that the circuit was optimized for. The horizontal solid and dashed lines indicate the DMRG estimates for the S0 and S1 eigenstates.
The active space considered for the [2Fe-2S] cluster is small enough for HCI to be able to find the relevant bitstrings by exploring and performing diagonalizations in subspaces significantly larger than the optimal ones. This shows that in some problem instances, the diagonalization itself is not the runtime bottleneck. Instead, the bottleneck is a sub-optimal proposal of electronic configurations. Larger and more strongly correlated systems will pose more challenges for the classical heuristics. It is these systems where quantum circuits can play a role.
V Experimental results
@@
V.1 Visualization of the wave function amplitudes
@@ V.2 Scaling and runtime
Fig. S12 shows the potential energy surfaces obtained by the estimator with different values of $d$, the number of determinants used for the subspace expansion and diagonalization. Increasing the value of $d$ improves the results both qualitatively and quantitatively. For the smaller values of $d$, the potential energy surface shows unphysical oscillations when the bond length is large. For larger values of $d$, not only the energy values are lower, but also the oscillations in the energy as a function of the bond length are significantly decreased in amplitude. Improvement of the ground state approximation with an increasing amount of computational resources is also observed in the energy-variance analyses of the main text (Fig. 4), and the supplementary materials (Figs. S21 and S22).
@@
V.3 Effect of orbital optimizations in the dissociation of N2 (cc-pVDZ)
@@
V.4 Orbital occupation numbers for the low-energy spectrum of [2Fe-2S]
@@
V.5 Quantum signal in the experiments
@@
The third set of tests involves comparing the performance of the HPC quantum estimator with configuration recovery run on configurations obtained from the quantum device and samples obtained from the uniform distribution. We consider the uniform distribution over all possible configurations in the Fock space, as well as the uniform distribution over configurations with the correct particle number, as in Fig. 4 (c) in the main text  [^1] When considering samples from the uniform distribution over electronic configurations on the correct particle sector, the configuration recovery method cannot be applied since it only recovers configurations with the wrong particle number.. The results for these last two sets of tests are shown in the coming subsections for the 40-, 52- and 72-qubit experiments.
V.5.1 N2 experiments